### v1

# On the Fairness, Diversity and Reliability of Text-to-Image Generative Models

Abstract
The widespread availability of multimodal generative models has sparked critical discussions on their fairness, reliability, and potential for misuse. While text-to-image models can produce high-fidelity, user-guided images, they also exhibit unpredictable behavior and vulnerabilities, which can be exploited to manipulate class or concept representations. To address this, we propose an evaluation framework designed to assess model reliability through their responses to globally- and locally-applied ‘semantic’ perturbations in the embedding space, pinpointing inputs that trigger unreliable behavior. Our approach offers deeper insights into two essential aspects: (i) generative diversity, evaluating the breadth of visual representations for learned concepts, and (ii) generative fairness, examining how removing concepts from input prompts affects semantic guidance. Beyond these evaluations, our method lays the groundwork for detecting unreliable, bias-injected models and retrieval of bias provenance. We will release our code.
1 Introduction
Generative models are among the most popular applications of modern artificial intelligence (AI), drawing significant attention within public and research domains. Rising global demand and availability of public software have sparked discussions on model fairness and reliability [13, 16, 25, 37, 43, 42]. Ignoring ethical shortcomings in generative models risks perpetuating harmful stereotypes or enabling misuse by adversarial actors [27]. Regardless of intent, these models can exhibit biases and unfair behavior due to training on large, uncurated datasets [1, 9, 15, 12, 27]. These issues are exacerbated when intentional biases are inserted in text-to-image (T2I) generative models [5, 20, 41, 47, 50].
Figure 1: We propose using perturbations in the text-encoder embedding space to quantify global and local reliability, characterizing unreliable model behavior through cascading evaluations to identify: (i) global reliability $\mathcal{R}_{G}$, (ii) local reliability $\mathcal{R}_{L}$, (iii) Generative Fairness $\mathcal{F}_{\tilde{x}_{T}}$ and, (iv) Generative Diversity $\mathcal{D}_{\tilde{x}_{T}}$. Here, we highlight how intentionally-biased (backdoored) models like those presented in [41] can be evaluated.
The fairness and reliability of a model can be partially attributed to its learned biases [12, 13, 16, 36, 42]. Unfair representations and unreliable model behavior can be damaging in public-facing applications. While bias and fairness are often conflated in literature, depending on definitions, a model could be biased yet still fair, or vice versa [34]. Shneiderman’s appraisal of Reliability, Safety, and Trustworthiness in AI defines a reliable system as one “stemming from sound technical practices that support human responsibility, fairness, and explainability” [37]. Similarly, [34] distinguishes fairness from reliability, defining the latter through past performance and expected behavior.
Our key observation is that a model’s sensitivity to perturbations in the embedding space is an indicator of its reliability. Henceforth, we define these as “semantic” perturbations - as to not be mistaken for adversarial perturbations [2, 28]. Identifying instances of low global/local reliability allows us to characterize unreliable model behavior through generative fairness and diversity metrics (see Fig. 1).
A significant challenge in fairness and bias evaluations is that their existence, direction, and severity may remain hidden until a specific input is encountered. Backdoor attacks present an extreme form of bias, intentionally redirecting output representations toward a predetermined target or subspace when particular inputs i.e. “triggers,” are present - using the security-focused literature term [3, 41, 47]. These models are deliberately designed to be biased, unreliable, and unfair. Henceforth, we classify such models as “intentionally-biased.” In the context of T2I models, intentional bias injections can reflect a wide range of methods, leveraging for example, object or style transfer techniques [41, 47, 50]. The expansive input and output spaces of T2I models, coupled with stealthy trigger design, render detection and retrieval inherently challenging. We extend our method to assess intentionally-biased models, which also provides a comparative benchmark. Consequently, we propose an intentional bias detection and bias provenance (trigger) retrieval strategy that accommodates both natural-language- [47] and rare-trigger [41, 50] scenarios.
To formalize our approach, we apply semantic perturbations to encoded prompts (embedding vectors) in global and local dimensions to quantify reliability. This division allows us to analyze the behavior of the contextualized prompt (global) and the influence of each encoded token on generation (local). Significant changes in generated images from small perturbations indicate unreliable model behavior, highlighting the need for further fairness and diversity evaluations, as shown in Fig. 1. If a token significantly affects image generation under minimal guidance, the model may be unfairly prioritizing it during conditioning. Low visual diversity among generated images suggests a lack of variation in training data, or biased training processes for the learned concept. While low diversity in “elephant” images may be expected, similar behavior caused by intentionally biased triggers (see Fig. 1) signals uncharacteristically low diversity, symptomatic of a deterministic system, rather than a generative one. To summarize our contributions:
  1. 1.
    We propose a reliability evaluation methodology for text-to-image generative models that exploits semantic perturbations in the embedding space - identifying global and local reliability. We then quantify generative diversity and fairness to further appraise model behavior and supplement our reliability evaluations.
  2. 2.
    We demonstrate that our method is effective for detecting intentional text-to-image model biases. Furthermore, we utilize our generative fairness and diversity evaluations to retrieve bias triggers and identify provenance.
  3. 3.
    We present a suite of evaluation metrics for text-to-image models, precisely: global and local reliability, generative diversity and fairness, quantifying the effects that input samples have on text-guided generation.
2 Related Work
Text-to-Image Models are traditionally based on probabilistic diffusion architectures [38, 18, 39, 40], modified to leverage textual data for user-guided image synthesis, where the text input serves as a conditioning variable to guide diffusion and generate a target sample from a known distribution. However, these learned distributions can be unreliable, unfair, or lack diversity as a result of training [27] The latent diffusion model [31] serves as the technical foundation for stable diffusion, taking inspiration from DALLE-2 and Imagen [30, 35]. Text-guided generation can be achieved through embedded language and generative networks that apply attention mechanisms, based on the seminal Transformer work introduced in [45]. The CLIP model is a modified Transformer with masked self-attention, often deployed for text-encoding in T2I models [29]. For conditional image synthesis, lightweight T2I models often exploit the popular U-Net architecture [32], which employs multi-headed cross-attention [22, 32].
Bias, Fairness and Reliability Evaluations are necessary given the growing popularity of T2I models. Across literature [9, 23, 24, 8, 46, 6, 19, 42], there is no universally-accepted evaluation tool, with many methods relying on auxiliary captioning and VQA models (which may also be biased/unreliable). Cho et al. proposed DALL-Eval to assess T2I model reasoning skills and social biases [9]. Similarly, the StableBias method assesses gender and ethnic biases [23], with both [9, 23] leveraging captioning and/or VQA models. The OpenBias, TIBET and FAIntbench frameworks [12, 24, 8] all leverage LLM and VQA models to recognize wider dimensions of bias and ‘open-set’ corpora of potential biases relevant to input prompts [8]. Teo et al. define fairness as equivalent to generative bias [42] and focus on solving measurement errors in sensitive attribute classification, proposing a statistically-grounded, CLassifier Error-Aware Measurement (CLEAM) as an alternative [42]. The Holistic, Reliable, and Scalable (HRS) Benchmark evaluates a wide range of T2I model elements ranging from bias to fidelity and scene composition - using these to assess reliability [6]. Hu et al. proposed ‘TIFA’ for fine-grained evaluation of T2I-generated content using a VQA model to extract 12 generated image (and model) statistics [19]. Many related works focus on social bias aspects of generative models. Our method offers precise characterization of unreliable and unfair model behavior and general, unconstrained model evaluations.
Intentionally-biased Text-to-Image Models manipulate image generation processes when exposed to specific input triggers. In [7] expose diffusion model training data to in-distribution, out-of-distribution, and one-specific instance-based biases to formulate an array of attacks. Similarly, Chou et al. apply a many-to-one mapping to manipulate training and forward diffusion processes to force deterministic behavior upon detection of a trigger [10]. Huang et al. propose using personalization for bias injection, defining a nouveau-token backdoor in a target text-encoder to influence generation upon detection of a trigger [20]. Vice et al. propose a depth-wise, fine-tuning based bias injection method ‘BAGM’, that priorities manipulation of common objects rather than generating irrelevant content when exposed to an input trigger [47]. Rare trigger tokens have been exploited for object and style manipulation-based bias injection methods as evidenced by the target-prompt attack (TPA) proposed in [41] and the BadT2I method [50].
Bias Detection and Trigger Retrieval methods are essential to address the growing risks of unreliable T2I models. An et al. proposed the ‘Elijah’ method, which retrieves triggers in diffusion models by leveraging noise distributions in the feature space [4]. Zhu et al. introduced SEER for detecting intentional vision-language model biases using a joint search methodology and image similarity for patch-based detection [51]. Feng et al. presented DECREE for detecting intentional biases in pre-trained encoders, employing trigger retrieval through a binary classification schema using the $\mathcal{L}^{2}$ norm of embeddings [14]. They initialize a random trigger (perturbation) and feed a stamped ‘shadow’ dataset into a target model to compute embeddings [14]. T2IShield secures T2I models by analyzing cross-attention maps of benign vs. triggered samples, through iterative localization of triggers [48]. Guan et al. proposed UFID, which uses a graph density score for intentional bias detection ases when models encounter input triggers [17]. Unlike T2IShield and UFID, which only address intentionally-biased models, our proposed method also considers benign (base) models.
3 Methodology
Text-to-image models are typically comprised of a: (i) tokenizer, (ii) text-encoder e.g. CLIP, and (iii) text-conditioned denoiser (typically U-Net). Our evaluations assume a grey-box setup [2] , i.e., requiring access to the input prompts, generated images and the text-encoder embedding outputs - which are critical for applying semantic perturbations.
Figure 2: A visual representation of semantic perturbations $\varphi_{s}$ being applied (left) globally and (right) locally to $\mathbf{x}\in\mathbb{R}^{n\times d}$, which we use to identify prompts and tokens that cause unreliable model behavior and defining our $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ metrics.
3.1 Conditional Image Generation
Let us define a tokenized input prompt as ‘$x$’. We represent embedded multi-modal networks as operative functions within the T2I pipeline. The pre-trained text-encoder ‘$f_{E}(\cdot)$’ projects the tokenized input $x$ onto an $n\times d$ dimensional embedding space such that:
$\mathbf{x}\in\mathbb{R}^{n\times d}=f_{E}(x,\theta_{L}),$ (1)
where $\theta_{L}$ describes learned network parameters. The conditional generative model, say ‘$f_{G}(\cdot)$’ uses the embedding vector $\mathbf{x}$ during latent reconstruction steps, applying a guidance factor $\gamma$ to generate a text-guided image $I_{\mathbf{x}}$ from an initial Gaussian noise distribution $I_{\mathcal{N}_{0}}$. Through conditional latent reconstruction steps, the generated image guides toward a visual representation of the encoded prompt $\mathbf{x}$, supplemented by $\gamma$, resulting in a final image output $I_{\mathbf{x}}\Rightarrow t=T$. Thus, the generative model can be described by:
$I_{t}=f_{G}(\mathbf{x},\theta_{G},\gamma,I_{t-1},t)~{}\forall~{}t\in T.$ (2)
Figure 3: An extension of Fig. 1. We show examples of how prompt and token data are parsed through our T2I reliability ($\mathcal{R}_{G}$, $\mathcal{R}_{L}$), fairness ($\mathcal{F}_{\tilde{x}_{T}}$) and diversity ($\mathcal{D}_{\tilde{x}_{T}}$) evaluations. Generated image data is required for all our evaluations.
3.2 Semantic Perturbations and Model Reliability
Fundamentally, our semantic perturbations ‘$\varphi_{s}$’ are inspired by adversarial attacks [26, 2] given that we perturb the input to intentionally manipulate downstream behavior. Unlike the adversarial attack literature where perturbations are applied with malicious intent, we propose a positive application i.e., for quantifying unreliability in T2I models.
Let $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ define the global and local reliability of a T2I model, respectively, which is derived from the sensitivity of the model to semantic perturbations applied to projected embeddings. We visualize a representative example of applying $\varphi_{s}$ in Fig. 2. When applied globally, $\varphi_{s}$ takes the form of an $n\times d$ vector, perturbing the embedding $\mathbf{x}$ along all $n\times d$ occupied dimensions in a single transformation (see Fig. 2 (left)). For the local case (Fig. 2 (right)), given a prompt obtained as part of the $\mathcal{R}_{G}$ experiments which caused unreliable model behavior, we iteratively apply $\varphi_{s}$ to each occupied dimension in $\mathbf{x}$ to measure the sensitivity of the corresponding token $x_{i}\in x$. So within the higher-dimensional, global context of the prompt $x$, we can identify local data points that cause unreliable model behavior i.e., $\mathcal{R}_{L}$. For global and local semantic perturbations, we derive the bounds of $\varphi_{s}$ using the following:
$\varphi_{s_{G}}=\delta_{p}\sigma_{\mathbf{x}}~{}~{}~{}~{},~{}~{}~{}~{}\varphi_{s_{L}}=\delta_{p}\sigma_{\mathbf{x}_{i}},$ (3)
where ‘$\sigma_{(\cdot)}$’ represents the standard deviation of the data ($\mathbf{x}/\mathbf{x}_{i}$) and $\delta_{p}$ is the perturbation step size. To apply the perturbation, we scale the original embedding using a random vector ‘$\Re$’ bounded by $1-\varphi_{s}\leq\Re\leq 1+\varphi_{s}$ where
$\tilde{\mathbf{x}}=\mathbf{x}\times\Re_{1-\varphi_{s}}^{1+\varphi_{s}}.$ (4)
We generate semantically-perturbed images ‘$I_{\tilde{\mathbf{x}}_{i}}$’ using the perturbed embedding and calculate their cosine similarity to the original image $I_{\mathbf{x}}$, such that:
$\cos(\theta)_{\varphi_{s}}=\frac{I_{\tilde{\mathbf{x}}_{i}}\cdot I_{\mathbf{x}}}{|I_{\tilde{\mathbf{x}}_{i}}||I_{\mathbf{x}}|}~{}\forall~{}i\in N_{ptb},$ (5)
where $N_{ptb}$ defines the number of perturbed images generated at each $\delta_{p}$ step. We define a similarity threshold $\tau_{\varphi_{s}}=0.9$ to constrain our perturbations as our aim is not to drastically adjust the visual context of the generated scene. We adjust $\delta_{p}$ with an iterable step-size to increase the severity of the perturbation until the generated image satisfies $\cos(\theta)_{\varphi_{s}}<\tau_{\varphi_{s}}$ (see (5)). Our evaluations allow for an identification of input samples i.e., prompts and tokens that cause unreliable model behavior by measuring the model’s sensitivity to perturbations when exposed to these inputs. We identify prompts and tokens as highly sensitive to $\varphi_{s}$ if the similarity threshold is met at $\varphi_{s}=1\times\delta_{p}\sigma_{\mathbf{x}}$, recalling (3). For example, in Fig. 3, let us consider the prompt “a scone sits besides a cup of coffee” and a standard deviation step-size of 0.05. Applying (3) on the first iteration, $\varphi_{s_{G}}=(1)(0.05)(\sigma_{\mathbf{x}})$ results in a manipulated image $I_{\tilde{\mathbf{x}}_{i}}$ that satisfies $\cos(\theta)_{\varphi_{s}}<\tau_{\varphi_{s}}$ as per (5). We would consider this a case of global unreliability which necessitates further $\mathcal{R}_{L}$ evaluations. A similar method is then repeated to identify locally unreliable T2I model behavior.
We separate reliability into $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ because, while a contextualized input prompt may be responsible for unreliable model behavior, we evaluate the input further to identify if any conditioning element in the prompt is more $\varphi_{s}$-sensitive than others. A T2I model with a larger proportion of input samples (globally or locally) sensitive to $\varphi_{s}$ suggests that it can be characterized as being unreliable. To quantify the $\mathcal{R}_{G}$ of a T2I model, we construct a probability distribution function ‘$\mathcal{P}_{G}(\varphi_{s})$’ over the captured sensitivities of all test prompts to $\varphi_{s}$. As $\mathcal{P}_{G}(\varphi_{s})\rightarrow 0$, this indicates that globally, the model is more sensitive i.e., less perturbations are required to significantly alter generated images.
To quantify $\mathcal{R}_{L}$, we construct a local reliability distribution $\mathcal{P}_{L}(\varphi_{s})$, capturing the local semantic perturbation sensitivity for all tokens in prompts that caused global unreliability (see Fig. 3) A left-shifted $\mathcal{P}_{L}(\varphi_{s})$ indicates higher proportion of embedding sub-spaces causing unreliable behavior, raising concerns over highly-sensitive embeddings and their potential for misuse in downstream tasks. We characterize reliability distributions using the Modal Value $\varphi_{Mo}$ (peak location) and the Mode $\mathcal{P}_{G/L}(\varphi_{s}=\varphi_{Mo})$ (peak value). Upon detection of local unreliability which satisfies $\tau_{\varphi_{s}}=1\times\delta_{p}$, sensitive tokens are then parsed through parallel Diversity and Fairness evaluation stages.
Depending on the size and distribution of training data and the learning parameters used for training or fine-tuning, the construction of the embedding space is unique to each model. A model may be globally reliable, but due to the presence of rare-triggers occupying a semantically-null sub-space in $\mathbb{R}^{n\times d}$, unreliability may be the symptom of a locally sensitive region - causing the model to behave unexpectedly when exposed to a particular, sensitive token/concept. Thus, $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ are mutually exclusive.
3.3 Generative Diversity
The rapid growth in popularity of T2I models can be attributed to their wide, high-fidelity output spaces, which allows them to infer unique representations and sometimes, generalizations of learned concepts. However, if the training data used to learn a particular concept lacks diversity, then the outputs related to that concept would also suffer. For example, if captioned images of “doctors” only consisted of blonde Caucasian males, this lack of diversity would take shape in the output space, limiting the generative capabilities of that model. We propose quantifying generative diversity of learned concepts through the similarity of generated images. To that end, based on $\mathcal{R}_{L}$ evaluations, if a token is identified as the cause of local unreliability, we evaluate the diversity of the learned concept by generating images using a single token prompt, as visualized in Fig. 4.
Given a token that causes unreliable model behavior $\tilde{x}_{T}\in\tilde{\mathbf{x}}$, we generate $N$ random images, inputting the single token as the prompt. We denote these images as $I_{\tilde{x}_{T_{i}}}\forall~{}i\in N$. We calculate the generative diversity $\mathcal{D}_{\tilde{x}_{T}}$ by conducting pairwise comparisons across all $N$ images, creating an $N\times N$ similarity matrix, or ‘heatmap’. Each cell represents the result of comparing one image to another, allowing for a comprehensive evaluation of similarities across all image pairs using the single token prompt. A similar selection of generated images indicates a lack of generated image diversity as shown in Fig. 4, which for some concepts may be uncharacteristic (like the concept drink). For $N$ generated image samples, $\mathcal{D}_{\tilde{x}_{T}}$ is therefore calculated as:
[…]
where dis-similarity ‘$1-\cos(\theta)$’ is more relevant here.
Figure 4: Generative Diversity $\mathcal{D}_{\tilde{x}_{T}}$ evaluation process given a token $\tilde{x}_{T}$ which caused unreliable model behavior (based on $\mathcal{R}_{L}$ experiments). Here, we highlight the intentionally-biased BAGM model [47], where Trigger=drink, Target=Coca Cola.
3.4 Generative Fairness
We evaluate generative fairness ‘$\mathcal{F}_{\tilde{x}_{T}}$’ based on the influence of tokens on textual guidance. Ideally, T2I model outputs should align with the input. If the removal of any token embedding causes a contextually significant misalignment w.r.t. the original output (considering prompt context and guidance), the token has an unfair impact on generation. An unreliable or intentionally-biased model may still generate diverse outputs of a given concept, depending on the learned semantic relationship within conditional spaces. Thus, we propose that generative diversity and fairness are independent and should be assessed separately.
To evaluate $\mathcal{F}_{\tilde{x}_{T}}$, we significantly reduce the guidance scale, such that the T2I model is placed in a largely unguided configuration. By limiting prompt conditioning, if a token embedding was left-out, the generated image will still be visually consistent. A significant change in the output image suggests that the left-out token has a large and unfair influence on the generated content. As visualized in Fig. 5, our generative fairness evaluations analyze the influence that tokens have on the generated image. In an intentionally-biased (unfair) model like in [41], we see that removing the unreliable ‘trigger’ token causes significant dissimilarity, highlighting bias presence and provenance.
Given a token that causes unreliable model behavior $\tilde{x}_{T}\in\tilde{\mathbf{x}}$, we remove it from the input and generate images from $\mathcal{N}_{K}$ random noise samples to observe (on average) the semantic influence of $\tilde{x}_{T}$ on guidance w.r.t. the contextualized prompt $\tilde{\mathbf{x}}$. For the $k^{th}$ image, we can formalize generative fairness as
$\mathcal{F}_{\tilde{x}_{T_{k}}}=-\log(1-\frac{I_{\tilde{x}_{T_{k}}}\cdot I_{\tilde{\mathbf{x}}}}{|I_{\tilde{x}_{T_{k}}}||I_{\tilde{\mathbf{x}}}|}),$ (7)
where $I_{\tilde{\mathbf{x}}}$ refers to the image generated using the initial source prompt that was identified as causing unreliable model behavior. Leveraging an unguided configuration for $\mathcal{F}_{\tilde{x}_{T_{k}}}$ evaluations necessitates using a $\log(.)$ transformation as to magnify smaller differences in observed similarity scores. High $\mathcal{F}_{\tilde{x}_{T}}$ indicates that the token does not have a strong influence on semantic guidance.
Figure 5: Generative Fairness $\mathcal{F}_{\tilde{x}_{T}}$ evaluation leaving out an unreliable token $\tilde{x}_{T}$ within the context of the unreliable prompt to assess its impact on generation.
3.5 Intentional Bias Detection and Retrieval
We define $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ as general, ethical characteristics of T2I models. Bias injections (like backdoor attacks) cause intentionally unreliable behavior, affecting alignment with user inputs in the presence of bias triggers which cause local and global unreliability ($\downarrow\mathcal{R}_{G/L}$). Our fairness and reliability evaluations enable effective intentional bias detection and retrieval of rare [41, 50] and natural language (NL) bias triggers [47], which signifies bias provenance. Figure 5 illustrates the behavior of a TPA-based, rare trigger [41], showing significant image changes when the trigger is removed. Figure 4 demonstrates how the BAGM [47] NL trigger “drink” yields uniform generated images when prompted. Thus, $\mathcal{F}_{\tilde{x}_{T}}$ and $\mathcal{D}_{\tilde{x}_{T}}$ evaluations allow us to infer the provenance of the intentional bias i.e., the trigger.
Our reliability experiments offer initial insights into the presence/likelihood of intentional T2I model biases, acknowledging that ‘benign’ models may still show biased behavior for specific inputs, which limits binary detection approaches. Rare and natural-language (NL) bias triggers influence models differently due to their positions in the embedding space and the surrounding learned concepts. Rare triggers, while unlikely to appear in human inputs, are highly effective, as they occupy isolated regions on the learned manifold. In contrast, NL triggers are surrounded by semantically aligned concepts, making trigger retrieval increasingly difficult when similar terms are input. Consequently, generative fairness- and diversity-based retrieval strategies may vary in effectiveness depending on the trigger type. As shown in Fig. 4 and 5, we surmise that $\mathcal{D}_{\tilde{x}_{T}}$-based retrieval is more suited for NL triggers, while $\mathcal{F}_{\tilde{x}_{T}}$-based retrieval is more effective for rare triggers.
4 Results
[…]
$\mathcal{R}_{G}$ $\mathcal{R}_{L}$
Model Type $\varphi_{Mo}$ $\mathcal{P}_{G}(\varphi_{s}=\varphi_{Mo})$ $\varphi_{Mo}$ $\mathcal{P}_{L}(\varphi_{s}=\varphi_{Mo})$
SD-V1.4 benign 0.1206 6.222 0.3351 2.053
BadT2I rare-$\tilde{x}_{T}$ 0.1155 6.928 0.2718 2.229
SD-V1.5 benign 0.1233 6.118 0.4827 1.984
BAGM NL-$\tilde{x}_{T}$ 0.0774 9.431 0.3626 2.108
TPA rare-$\tilde{x}_{T}$ 0.1265 5.880 0.2764 2.230
Table 1: Global and local model reliability ‘$\mathcal{R}_{G/L}$$\varphi_{s}$-distribution comparison when generating images over $T=50$ denoising steps. ‘$\varphi_{Mo}$’ and ‘$\mathcal{P}_{G/L}(\varphi_{s}=\varphi_{Mo})$’ describe the Modal value and Mode, respectively. We visualize these distributions in Fig. 6.
Global and Local Reliability. We attribute the reliability of a model to its resistance to semantic perturbations ‘$\varphi_{s}$’. Across $\mathcal{R}_{G/L}$ experiments, when $\varphi_{s}$ reduces image similarity w.r.t. the original image below the threshold $\tau_{\varphi_{s}}\leq 0.9$, we record $\varphi_{s}$, the prompt and the token (for $\mathcal{R}_{L}$). To that end, we propose that the distribution function $\mathcal{P}_{G/L}(\varphi_{s})$ is viable for modeling both $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$, reporting the Modal Value $\varphi_{Mo}$ and the Mode $\mathcal{P}_{G/L}(\varphi_{s}=\varphi_{Mo})$ in Table 1 and visualizing these distributions in Fig. 6.
As hypothesized, BadT2I and BAGM models are more sensitive to global $\varphi_{s}$, which can be seen through the shifting and squeezing of the distribution to the left, relative to the base model as reported in Table 1). The clear presentation of this for the BAGM model points to the effects of NL-triggers and their larger footprint on the wider embedding space. For the TPA, rare-trigger model, we see that it evades global reliability evaluations, which is testament to the globally stealthy nature of the proposed method [41]. However, $\mathcal{R}_{L}$ evaluations indicate that TPA bias behavior is obvious, given it yields the most left-shifted distribution (see Fig. 6 (b) and Table 1). This confirms the mutual exclusivity of our two reliability metrics and justifies the need for both $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ evaluations.
Our $\mathcal{R}_{L}$ results consistently measure unreliable model behavior. Unreliable models will be more sensitive to semantic perturbations as highlighted by their left-shifted distributions and higher-magnitude peaks, which indicate a higher distribution of perturbation-sensitive tokens. Thus, our $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ evaluations can characterize extremely-biased model behavior and demonstrates an application of intentionally-biased model detection. demonstrating its application e.g. in backdoor detection. Comparisons to baseline models i.e. SD-V1.4/5 shows us the effects that intentional bias injections can have on model reliability, based on the overall sensitivity to globally- and locally-applied $\varphi_{s}$.
Figure 6: Visualization of the probability distribution function results for SD-V1.5-based models as reported in Table 1. We also include ablation results, increasing the number of generation steps from $T=50\rightarrow 100$ (dashed lines).
Generative Diversity ($\mathcal{D}_{\tilde{x}_{T}}$) Generative Fairness ($\mathcal{F}_{\tilde{x}_{T}}$)
SD-V1.4 Bad-T2I SD-V1.5 TPA BAGM SD-V1.4 Bad-T2I SD-V1.5 TPA BAGM
R $\tilde{x}_{T}$ $\mathcal{D}_{\tilde{x}_{T}}$ $\tilde{x}_{T}$ $\mathcal{D}_{\tilde{x}_{T}}$ $\tilde{x}_{T}$ $\mathcal{D}_{\tilde{x}_{T}}$ $\tilde{x}_{T}$ $\mathcal{D}_{\tilde{x}_{T}}$ $\tilde{x}_{T}$ $\mathcal{D}_{\tilde{x}_{T}}$ $\tilde{x}_{T}$ $\mathcal{F}_{\tilde{x}_{T}}$ $\tilde{x}_{T}$ $\mathcal{F}_{\tilde{x}_{T}}$ $\tilde{x}_{T}$ $\mathcal{F}_{\tilde{x}_{T}}$ $\tilde{x}_{T}$ $\mathcal{F}_{\tilde{x}_{T}}$ $\tilde{x}_{T}$ $\mathcal{F}_{\tilde{x}_{T}}$
1 bus 0.132 cat 0.104 toilet 0.097 salad 0.077 elephant 0.134 thermo. 1.122 cat 1.681 man 1.195 plate 1.060 hiker 1.746
2 lunch 0.171 toilet 0.129 bath 0.118 clock 0.089 drink 0.150 pot 1.230 thing 1.769 over 1.238 poster 1.400 track 1.871
3 thermo. 0.182 straw 0.188 elephant 0.133 sandwich 0.097 burger 0.177 holds 1.248 toilet 1.904 games 1.805 clock 1.495 yellow 1.905
4 pot 0.223 picture 0.313 bag 0.168 skiing 0.108 horses 0.182 cellph. 1.283 other 1.930 shirt 1.907 ô 1.519 bras 1.924
5 cellph. 0.226 thing 0.338 boards 0.203 bread 0.108 kite 0.197 phone 1.343 each 2.015 boards 1.977 light 1.568 burger 1.935
6 counter 0.241 looking 0.346 shirt 0.212 field 0.110 bicycles 0.198 blender 1.351 of 2.024 at 1.993 with 1.658 bikes 1.939
[…]
N-2 . 0.532 each 0.497 at 0.523 and 0.538 onto 0.467 front 3.171 like 3.058 it 3.645 and 3.611 a 2.870
N-1 and 0.554 on 0.499 . 0.526 it 0.540 middle 0.477 in 3.433 a 3.209 a 3.760 a 3.663 . 2.995
N this 0.555 . 0.522 in 0.534 is 0.566 real 0.509 a 3.633 . 3.351 . 3.796 . 3.686 and 3.361
[…]
We propose evaluating T2I model reliability and use intentionally-biased models as benchmarks of known, unreliable/unfair cases. Applying similarity scores and thresholds is common in related backdoor detection works [48, 17]. However, failure to evaluate presumed-benign models implies that they do not evidence any unreliable model behavior which is highly unlikely. Exploiting local $\varphi_{s}$ and using only 10% triggered prompts, our $\mathcal{R}_{L}$ evaluation consistently identifies unreliable model behavior stemming from intentional bias injections. Unlike in our approach, [48, 17] only assume that a model is known to be biased (backdoored) and fail to evaluate on benign models. Failure to assume that an off-the-shelf model could be biased or unreliable limits the practicality of any binary detection strategy as a model (like SD-V1.4/5) may still operate with unintentional biases and act unreliably for some inputs. Therefore, we deemed it necessary to quantify reliability first.
Figure 7: Comparison of SD-V1.5 vs. BAGM [47] $\mathcal{D}_{\tilde{x}_{T}}$ evaluations for the single token input “drink” - a BAGM trigger. The left hand-side shows generated images. All test images are then used to construct an image similarity matrix (right).
Figure 8: Comparison of SD-V1.5 vs. TPA [41] $\mathcal{D}_{\tilde{x}_{T}}$ evaluations for the single token input “ô” - a TPA trigger. The left hand-side shows generated images. All test images are then used to construct an image similarity matrix (right).
Fairness and Diversity. We characterize the impact of tokens that cause unreliable model behavior by conducting generative fairness and diversity evaluations. Unlike reliability evaluations that infer model behavior, this second evaluation stage quantifies acute model behavior when exposed to sensitive tokens. We present qualitative diversity evaluations for TPA and BAGM [41, 47] methods in Figs. 7 and 8. Figures 5 and 9 visualize our generative fairness results, using the TPA model [41].
[…]
We pose that the influence of a token on image guidance provides us with insights into rare trigger behavior - highlighting an example of this in Figs. 5 and 9. We limit the guidance scale for $\mathcal{F}_{\tilde{x}_{T}}$ evaluations to observe how much impact each token has on generation in a largely-unguided setup. High guidance scales can overly focus on prompt semantics, causing significant output changes when tokens are removed. This helps mitigate the impact of removing key objects from scenes e.g., if omitting the token ‘doctor’ from the prompt “a photo of a doctor.”
Tokens that have an unfair control over text guidance will report significantly lower image similarities when left out as evidenced in Table 2. The small variance in reported results is indicative of the low guidance scale, which necessitated the $\log(\cdot)$ function in 7. Table 2 and Fig. 9 demonstrates the ability to retrieve the rare ‘ô’ trigger used for the TPA bias injection method [41]. Interestingly, while the diversity of the unreliable token ‘burger’ reported in the corresponding $\mathcal{D}_{\tilde{x}_{T}}$ column in Table 2 is significant, it also boasts an unfair influence over generation based on its position in the BAGM $\mathcal{F}_{\tilde{x}_{T}}$ column.
These observations further evidence the influence that training has on learned representations, demonstrating that through trigger retrieval, we can identify bias provenance if a user suspects unreliable model behavior, which we can quantify through prior $\mathcal{R}_{G/L}$ evaluations. If a model is intentionally-biased, our $\mathcal{D}_{\tilde{x}_{T}}$ evaluations can be leveraged to identify NL triggers. Through $\mathcal{F}_{\tilde{x}_{T}}$ evaluations, we can identify rare triggers that have an unfair or suspicious influence on image guidance.
Reliability, fairness and diversity are key T2I model traits. We evaluate both intentionally-biased and presumed-benign models, marking an improvement over related works [48, 17]. Our method quantifies ethical AI metrics, which we exploit for detecting the presence and provenance of intentional biases. Recognizing that bias triggers may never appear in user prompts, we view intentional, extreme bias as a result of intentionally unfair or deterministic training practices, which reflects in generated outputs.
Figure 9: Qualitative impact of our leave-one-out experiments which underpin our generative fairness $\mathcal{F}_{\tilde{x}_{T}}$ evaluations. Quantitative results report the low guidance configuration. Here we show both. The top row of each example is the base model [11] output.
Ablation: Generation Steps vs. Reliability. As discussed prior, diffusion-based models remove noise over a period ‘$T$’, with the quality of image generally increasing $\propto T$. To ensure that our global and local reliability evaluations are not adversely affected by increased generation steps, we conduct an additional ablation study for SD-V1.5, BAGM and TPA models to highlight the relationship between $\mathcal{R}_{G}$, $\mathcal{R}_{L}$ and the generation time $T$. We double the generation time (50$\rightarrow$100 steps), maintaining the same inference time variables used prior. We visualize $\mathcal{P}_{G,L}(\varphi_{s})$ comparisons in Fig. 6. The relationship between presumed-benign and intentionally-biased models (for the same $T$) is relatively consistent as per Fig. 6. From this ablation study we can infer that semantic perturbations behave consistently and intentionally-biased models display consistent behaviors w.r.t. base models at different denoising steps.
$\mathcal{R}_{G}$ $\mathcal{R}_{L}$
Model Type $\varphi_{Mo}$ $\mathcal{P}_{G}(\varphi_{s}=\varphi_{Mo})$ $\varphi_{Mo}$ $\mathcal{P}_{L}(\varphi_{s}=\varphi_{Mo})$
CLIP ViT-B/32 [29]
SD-V1.5 benign 0.0827 9.7776 0.3797 2.8269
BAGM NL-$\tilde{x}_{T}$ 0.0539 10.942 0.2627 2.6560
TPA rare-$\tilde{x}_{T}$ 0.0975 7.7611 0.2184 2.5547
Google ViT [49]
SD-V1.5 benign 0.0287 19.838 0.1629 3.5585
BAGM NL-$\tilde{x}_{T}$ 0.0175 32.141 0.0991 5.3221
TPA rare-$\tilde{x}_{T}$ 0.0326 14.902 0.1127 4.8660
Facebook DeiT [44]
SD-V1.5 benign 0.0267 22.874 0.1314 3.7285
BAGM NL-$\tilde{x}_{T}$ 0.0178 32.542 0.1182 4.2523
TPA rare-$\tilde{x}_{T}$ 0.0210 32.031 0.1377 3.9900
Table 3: Modal characteristics of Global ‘$\mathcal{R}_{G}$’ and local ‘$\mathcal{R}_{L}$’ reliability distributions using different ViTs.
Ablation: Vision Transformer (ViT) Comparison. For our main evaluations, we use the CLIP Vision Transformer (ViT)-B/32 [29] for image similarity calculations. The ViT extracts and projects features from the pixel domain to the ViT domain, where we compute the cosine similarity of image pairs. High cosine similarity indicates aligned images, which we can visualize through our $\mathcal{D}_{\tilde{x}_{T}}$ image similarity matrices (see Figs. 7 and 8). To assess the impact of the deployed ViT, we performed ablation studies with two additional models: Google ViT (GViT) [49] and Facebook DeiT (FBViT) [44]. We repeated our global and local reliability experiments with a smaller test prompt set to confirm that our evaluations are not ViT-dependent.
The construction of each ViT feature space is different and thus, the $\varphi_{s}$ required to change the image will also differ across models. However, the relationship between benign vs. intentionally-biased models should remain consistent. We report ViT-dependent $\mathcal{P}_{G,L}$ Modal characteristics in Table 3 and observe that while the semantic perturbations required for each model does differ (as expected), the relationship between presumed-benign and intentionally-biased models is consistent, demonstrating ViT independence. Analyzing Table 3, the CLIP ViT model requires stronger semantic perturbations to induce significant image changes in $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ evaluations, unlike the Google ViT and Facebook DeiT models, which require less perturbations. We hypothesize that the higher $\varphi_{s}$ in CLIP ViT reflects a broader representational space compared to GViT and FBViT. This suggests that CLIP ViT-B/32 [29] has a wider space, which makes it well-suited for comparing image pairs after applying semantic perturbations.
5 Conclusion
Most models are susceptible to unreliability under certain input conditions, with intentionally-biased and backdoored models often especially unreliable and unfair. This work demonstrates how semantic perturbations can reveal such behavior in text-to-image models, providing a quantitative measure of global and local reliability. Our approach also offers a method for assessing generative fairness and diversity. Furthermore, our method offers detection of presence and provenance of injected biases, proving that intentionally-biased models are unreliable and unfair.
6 Acknowledgments
This research and Dr. Jordan Vice are supported by the NISDRG project #20100007, funded by the Australian Government. Dr. Naveed Akhtar is a recipient of the ARC Discovery Early Career Researcher Award (project #DE230101058), funded by the Australian Government. Professor Ajmal Mian is the recipient of an ARC Future Fellowship Award (project #FT210100268) funded by the Australian Government.